An aluminum block (c = 9.1 X 102 J/kg K) of mass 0.50 kg at a temperature of 155 degrees Celsius is dropped into an aluminum calorimeter cup of mass.200 kg containing 1.00 kg of water ( c = 4.2 X 103 J/kg K) at 20 degrees Celsius. The system is insulated and attains equilibrium at a final temperature Tf

(I) Using the definition of specific heat capacity, set up the Qtotal = 0 equation.

(II) solve for Tf


Calculate the temperature of a fluid when both Fahrenheit and a Celsius thermometer are immersed in it, under the following conditions (a) the numerical reading is identical in both thermometers, (b) the Fahrenheit is numerically twice that of the Celsius reading. Express the values in 0R and 0K.



1
Expert's answer
2021-04-14T11:31:47-0400

Solution.

"c_a = 9.1 \sdot 10^2 J/(kg K);"

"m_a=0.50kg;"

"t_a=155^oC;"

"m_c=0.200kg;"

"m_w=1.00kg;"

"c_w=4.2\sdot10^3J/(kgK);"

"t_w=20^oC;"

"T_f-?;"

"Q_a=c_am_a(T_a-T_f);"

"Q_c=c_cm_c(T_f-T_c);"

"Q_w=c_wm_w(T_f-T_w);"

"a) If Q_{total}=0, c_am_a(T_a-T_f)+c_cm_c(T_f-T_c)+c_wm_w(T_f-T_w)=0;"

"b)c_am_a(T_a-T_f)=c_cm_c(T_f-T_c)+c_wm_w(T_f-T_w)\implies"

"T_f=\dfrac{c_a m_a T_a+c_cm_cT_c+c_wm_wT_w}{c_am_a+c_cm_c+c_wm_w};"

"T_f=\dfrac{9.1\sdot10^2\sdot0.50\sdot155+4.2\sdot10^3\sdot1.00\sdot20+9.1\sdot10^2\sdot0.200\sdot20}{9.1\sdot10^2\sdot0.50+4.2\sdot10^3\sdot1.00+9.1\sdot10^2\sdot0.200}=33^oC;"

The Fahrenheit scale and the Celsius scale "intersect" at −40 ° (−40 ° F = −40 ° C).

"-40^oC=233^oK=419.67\u00baR;"

Fahrenheit (° F) = Celsius (° C) * 1,8 + 32 ;

"160^oC=320^oF=433^oK=779.67\u00baR;"

Answer: "T_f=33^oC;"

"-40^oC=233^oK=419.67\u00baR;"

"160^oC=320^oF=433^oK=779.67\u00baR;"